OpenStudy (anonymous):

(To convert the Celsius temperature C to the Fahrenheit temperature F) 5x −4y = 8, for y how do i write the answer in the form y=mx+b.

OpenStudy (anonymous):

isolate the y value: 4y=5x-8 y=(5/4)x-2

OpenStudy (anonymous):

but for my to get the answer i was told to do it with 4 steps

OpenStudy (anonymous):

me

OpenStudy (anonymous):

5x −4y = 8 4y=5x-8 the left and right side of the equation divide it by 4 y=(5/4)x-2

OpenStudy (anonymous):

and the next 3 step

OpenStudy (anonymous):

do anny body know the 3 step to (To convert the Celsius temperature C to the Fahrenheit temperature F) 5x −4y = 8, for y